Solve the given problem. Find the following and show your complete solution.


a. Critical Value


b. Margin of Error


c. Confidence Interval


GIVEN:


The average test scores in a Science class is normally distributed with a standard deviation of 8. 200 scores with a sample mean of 170 were selected at random. Estimate the population mean with:


a. 85% confidence


b. 90% confidence


c. 95% confidence


1
Expert's answer
2021-05-03T08:38:00-0400

a)85% confidence

i)The critical value is 1.44

ii)the margin of error= critical value*standard deviation

=1.44*8.200

=11.808

iii) confidence interval=mean±margin of error

=170±11.808

(158.192,181.808)

b)90% confidence

i) the critical value is 1.645

ii)the margin of error= critical value*standard deviation

=1.645*8.200

=13.489

iii) confidence interval=mean±margin of error

=170±13.489

(156.511,183.489)

c)95% confidence

i)The critical value is 1.96

ii)the margin of error= critical value*standard deviation

=1.96*8.200

=16.072

iii) confidence interval=mean±margin of error

=170±16.072

(153.928,186.072)